OpenStudy (anonymous):

solve for x . |x+1|-|x|+2|x-1|=2x-1

OpenStudy (anonymous):

we can solve this using by sign chart x<-1 -1 0 1 x+1 ---------0++++++++++++++++++++ x --------------------0++++++++++++++ x-1 ---------------------------------0+++++ then take the cases & equating with 2x-1 Case1 x<-1 _(x+1)+x-2(x-1)=2x-1 solve this u will get x=1/2 Case2(-1,0) x+1+x-2(x-1)=2x-1 x=2 Case3 (0,1) x+1-x-2(x-1)=2x-1 x=1 Case4 x>1 x+1-x+2(x-1)=2x-1 x=All real numbers Finally The solution sets are 1/2,2,1 Or u can check this values!!!1

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

thnx